# SAMba-UNet: Synergizing SAM2 and Mamba in UNet with Heterogeneous Aggregation for Cardiac MRI Segmentation

### v1

@@ 4.2 Evaluation Metrics
4.3 Quantitative Comparison
Table 1: Performance of Different Models on the ACDC Dataset
Model mDice↑ RV MYO LV
UNetRonneberger et al. (2015) 0.8993 0.8682 0.8776 0.9521
UNet++Zhou et al. (2019) 0.8994 0.8730 0.8740 0.9507
R50 Attn-UNetOktay et al. (2018) 0.8675 0.8758 0.7920 0.9347
VIT-CUPChen et al. (2021) 0.8145 0.8146 0.7071 0.9218
R50 ViT-CUPChen et al. (2021) 0.8757 0.8607 0.8188 0.9475
TransUNetChen et al. (2021) 0.8923 0.8591 0.8667 0.9511
SwinUNetCao et al. (2022) 0.8928 0.8701 0.8637 0.9449
UNETRHatamizadeh et al. (2022) 0.8861 0.8529 0.8652 0.9402
LeViT-UNet-384Xu et al. (2023) 0.9032 0.8955 0.8764 0.9376
Mamba-UNetWang et al. (2024) 0.8997 0.8817 0.8860 0.9313
SAMba-UNet(Our) 0.9103 0.9039 0.8935 0.9335
As demonstrated in the quantitative analysis of Table 1, SAMba-UNet achieves superior segmentation performance on the ACDC dataset. With an mDice score of 0.9103, it surpasses all comparative models by 0.71 percentage points over the suboptimal LeViT-UNet-384. Notably, the model exhibits performance gains of 0.94% and 0.84% for segmenting morphologically complex structures — the right ventricle (RV, 0.9039) and myocardium (MYO, 0.8935), respectively — which substantiates the effectiveness of its multi-scale feature modeling. Compared to conventional convolutional architectures (UNet++), attention-based methods (R50 Attn-UNet), and pure Transformer models (SwinUNet, UNETR), SAMba-UNet maintains comparable left ventricle (LV, 0.9335) segmentation accuracy (difference <2%) with mainstream approaches while achieving breakthrough collaborative segmentation performance through the novel integration of SAM2 architecture and state space models.
5 Ablation Study
We conducted experiments to explore the following two aspects: (1)The effectiveness of different components of the model architecture; (2)Investigation of Effectiveness in Different Adapters and Their Components.
5.1 The effectiveness of different components of the model architecture
As shown in the ablation study results in Table 2, the complete model architecture (ALL) achieves optimal performance across all evaluation metrics, fully demonstrating the importance of synergistic interactions between modules. Further analysis reveals that removing critical components (e.g., the OCA module) leads to significant degradation in segmentation accuracy and weakened boundary localization capability, underscoring its critical role in the system. While the removal of other components (such as IRMLP and AdaptAttn) does not completely compromise basic model functionality, all exhibit varying degrees of negative impacts on core performance metrics. These observations effectively validate the efficient design of the overall architecture and the functional complementarity among different modules.
Table 2: The effectiveness of different components of the model architecture.
@@ 5.1 The effectiveness of different components of the model architecture
5.2 Investigation of Effectiveness in Different Adapters and Their Components
As shown in the ablation study results of Table 3, the complete adapter architecture (ALL) achieves optimal performance across all evaluation metrics, fully demonstrating the effectiveness of the collaborative working mechanism among components. Removing core modules (such as the Channel Attention or Refiner) leads to significant performance degradation, particularly evident in boundary localization accuracy and segmentation consistency metrics. Although the absence of the MLP-Adapter has a relatively minor impact on overall performance, it still causes precision loss in detailed features, indicating the module’s irreplaceable role in feature refinement. Experimental results demonstrate that each submodule specifically enhances different capability dimensions of the model (e.g., region recognition accuracy, edge sharpness), while systematic integration of these purposefully designed components constitutes the key factor in achieving optimal segmentation performance.
Table 3: Investigation of Effectiveness in Different Adapters and Their Components.
@@ 5.2 Investigation of Effectiveness in Different Adapters and Their Components
6 Conclusion
The proposed SAMba-UNet innovatively integrates SAM2, Mamba, and UNet architectures, successfully addressing domain adaptation issues and fine-grained feature extraction challenges in medical image segmentation. Experimental results demonstrate that the Dynamic Feature Fusion Refiner effectively mitigates semantic discrepancies between natural and medical images, while the Heterogeneous Omni-Attention Convergence Module (HOACM) significantly enhances the collaborative modeling of global context and local details. On the ACDC dataset, the model outperforms existing methods in boundary segmentation accuracy for key cardiac structures (e.g., myocardium and ventricles) and sensitivity to complex pathologies.

### v2

@@ IV-C Evaluation Metrics
IV-D State-of-the-Art Comparison
As demonstrated in the quantitative analysis of Table I, SAMba-UNet achieves superior segmentation performance on the ACDC dataset. For models with officially released implementations, all experiments were conducted by re-training and testing on an A800 server (or using official weights directly), while for models without open-source implementations, we adopted the officially reported results from their respective papers. With an mDice score of 0.9103, it surpasses all comparative models by 0.71 percentage points over the suboptimal LeViT-UNet-384. Notably, the model exhibits performance gains of 0.94% and 0.84% for segmenting morphologically complex structures - the right ventricle (RV, 0.9039) and myocardium (MYO, 0.8935), respectively - which substantiates the effectiveness of its multi-scale feature modeling. Compared to conventional convolutional architectures (UNet++), attention-based methods (R50 Attn-UNet), and pure Transformer models (SwinUNet, UNETR), SAMba-UNet maintains comparable left ventricle (LV, 0.9335) segmentation accuracy (difference <<2%) with mainstream approaches while achieving breakthrough collaborative segmentation performance through the novel integration of SAM2 architecture and state space models. The qualitative visualization comparison results of different models are presented in Figure 5.
TABLE I: Performance of Different Models on the ACDC Dataset.
Method mDice↑ RV MYO LV
UNet [2] 0.8993 0.8682 0.8776 0.9521
UNet++ [3] 0.8994 0.8730 0.8740 0.9507
R50 Attn-UNet [38] 0.8675 0.8758 0.7920 0.9347
VIT-CUP [10] 0.8145 0.8146 0.7071 0.9218
R50 ViT-CUP [10] 0.8757 0.8607 0.8188 0.9475
TransUNet [10] 0.8923 0.8591 0.8667 0.9511
SwinUNet [39] 0.8928 0.8701 0.8637 0.9449
UNETR [40] 0.8861 0.8529 0.8652 0.9402
LeViT-UNet-384 [41] 0.9032 0.8955 0.8764 0.9376
Mamba-UNet [31] 0.8997 0.8817 0.8860 0.9313
SAMba-UNet (Ours) 0.9103 0.9039 0.8935 0.9335
Figure 5: Comparison of cardiac structure segmentation performance among different models on the ACDC dataset, where red boxes highlight regions with noticeable segmentation discrepancies between models.
TABLE II: The effectiveness of different components of the model architecture.
@@ IV-D State-of-the-Art Comparison
TABLE III: Investigation of Effectiveness in Different Adapters and Their Components.
@@
IV-E Ablation Study
We conducted experiments to explore the following two aspects.
Effectiveness of Different Components of Model Architecture. As shown in the ablation study results in Table II, the complete model architecture (ALL) achieves optimal performance across all evaluation metrics, fully demonstrating the importance of synergistic interactions between modules. Further analysis reveals that removing critical components (e.g., the OCA module) leads to significant degradation in segmentation accuracy and weakened boundary localization capability, underscoring its critical role in the system. While the removal of other components (such as IRMLP and AdaptAttn) does not completely compromise basic model functionality, all exhibit varying degrees of negative impacts on core performance metrics. These observations effectively validate the efficient design of the overall architecture and the functional complementarity among different modules.
Investigation of Effectiveness in Different Adapters and Their Components. As shown in the ablation study results of Table III, the complete adapter architecture (ALL) achieves optimal performance across all evaluation metrics, fully demonstrating the effectiveness of the collaborative working mechanism among components. Removing core modules (such as the Channel Attention or Refiner) leads to significant performance degradation, particularly evident in boundary localization accuracy and segmentation consistency metrics. Although the absence of the MLP-Adapter has a relatively minor impact on overall performance, it still causes precision loss in detailed features, indicating the module’s irreplaceable role in feature refinement. Experimental results demonstrate that each submodule specifically enhances different capability dimensions of the model (e.g., region recognition accuracy, edge sharpness), while the systematic integration of these purposefully designed components constitutes the key factor in achieving optimal segmentation performance.
V Conclusion
The proposed SAMba-UNet innovatively integrates SAM2, Mamba, and UNet architectures, successfully addressing domain adaptation issues and fine-grained feature extraction challenges in medical image segmentation. Experimental results demonstrate that the Dynamic Feature Fusion Refiner effectively mitigates semantic discrepancies between natural and medical images, while the Heterogeneous Omni-Attention Convergence Module (HOACM) significantly enhances the collaborative modeling of global context and local details. On the ACDC dataset, the model outperforms existing methods in boundary segmentation accuracy for key cardiac structures (e.g., myocardium and ventricles) and sensitivity to complex pathologies. These capabilities enable the model to serve as a high-precision perception module within an intelligent robot control system for computer-aided diagnosis and surgical assistance. It provides reliable multi-scale anatomical segmentation support for preoperative planning, real-time intraoperative navigation using augmented reality, and postoperative complication screening.